OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? 44 36 76 60

OpenStudy (nurali):

Let the number of women who auditioned be 'w'. Number of men is '100 - w'. Given, 2(100 - w) - 3w = 20 => 200 - 2w - 3w = 20 => 180 = 5w => w = 180/5 = 36 Hence, the number of women who auditioned is 36.